# Real-Time Dynamics in Two Dimensions with Tensor Network States via Time-Dependent Variational Monte Carlo

### v2

@@ IV.3 $\mathbb{Z}_{2}$ lattice gauge theory with Higgs field
Next, using the gauge-invariant PEPS–VMC framework recently introduced in Ref. [27], we examine the confinement dynamics in the $\mathbb{Z}_{2}$ gauge theory coupled to a Higgs field,
@@ IV.4 Superfluid bosons
Figure 5: Real-time density response of a 2D bosonic superfluid to a moving obstacle (the red dot). Snapshots show the deviation of the local density from 0.5 on a $13\times 13$ lattice for different times $T$ as the obstacle is dragged along the $x$ direction. (a) $|\bm{\mathbf{v}}|=0.5$. (b) $|\bm{\mathbf{v}}|=5.0$.
@@
where $\bm{\mathbf{R}}(t)$ is the position of the obstacle at time $t$. By Galilean invariance, we keep the superfulid at rest and drag the obstacle rightward at constant velocity: $\bm{\mathbf{R}}(t)=\bm{\mathbf{R}}_{0}+\bm{\mathbf{v}}t$. $H_{\text{obs}}(t)$ can also be realized via a laser in cold-atom simulations. The initial state is the ground state of $H_{\text{sf}}+H_{\text{obs}}(0)$. As shown in Fig. 5, for slow motion $|\bm{\mathbf{v}}|=0.5$, the density perturbation remains localized and approximately symmetric around the impurity, with no pronounced wake, consistent with dissipationless flow below the critical velocity. For fast motion $|\bm{\mathbf{v}}|=5.0$, the impurity leaves behind a pronounced density wake and radiates excitations throughout the system above the critical velocity. In particular, the density depletion caused by the obstacle closely follows the obstacle when $\bm{\mathbf{v}}$ is small, while it lags far behind when $\bm{\mathbf{v}}$ is large. This qualitative difference signals the prescence of a critical superfluid velocity, as anticipated by Landau [71].
@@ V Discussion and Outlook
The examples explored in Sec. IV demonstrate the breadth of physical phenomena accessible with this approach. These results establish PEPS–tVMC as a general-purpose tool for simulating elementary excitations, transport phenomena, and real-time response in a wide range of quantum systems with realistic computational resources. The improvements developed here apply equally to ground-state PEPS–VMC calculations [23, 24, 25, 26]. Extensions to excited states, finite temperature [73], and open quantum systems are also promising avenues. In the case of finite temperature, if purification-based ansatz is used, the small-$o$ trick can save the memory cost by a factor of 16 for spinful fermions!
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [74, 75]. The code for production is implemented with JAX [76]. Y.W. is supported by a startup grant from the IOP-CAS. I thank my colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with me. I benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Jannes Nys, Sheng-Hsuan Lin, Sajant Anand, Taige Wang, Michael Zaletel, Frank Pollmann, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu.
@@
    To implement this, one loops over $k$ and use an accumulator for the sum. This takes the same memory by a factor of $K$ with no additional time cost. ($oo^{\dagger}$ is $K$-times faster than $OO^{\dagger}$, but one executes $K$ of them in sequence. The memory cost of $p_{sa}$ is neglegible because its dtype is jnp.int8 compared to $O$’s dtype jnp.complex128).
@@ 5 OOdag : (Ns, Ns), obtained from get_OOdag()
  • dtdt: step of the time evolution

### v3

@@ IV.3 $\mathbb{Z}_{2}$ lattice gauge theory with Higgs field
Next, using the gauge-invariant PEPS–VMC framework recently introduced in Ref. [28], we examine the confinement dynamics in the $\mathbb{Z}_{2}$ lattice gauge theory coupled to a Higgs field,
@@ IV.4 Superfluid bosons
[…]
@@
where $\bm{\mathbf{R}}(t)$ is the position of the obstacle at time $t$. By Galilean invariance, we keep the superfluid at rest and drag the obstacle rightward at constant velocity: $\bm{\mathbf{R}}(t)=\bm{\mathbf{R}}_{0}+\bm{\mathbf{v}}t$. $H_{\text{obs}}(t)$ can also be realized via a laser in cold-atom simulations. The initial state is the ground state of $H_{\text{sf}}+H_{\text{obs}}(0)$. As shown in Fig. 5 (a,b), for slow motion $|\bm{\mathbf{v}}|=0.5$, the density perturbation remains localized and approximately symmetric around the impurity, with no pronounced wake, consistent with dissipationless flow below the critical velocity. For fast motion $|\bm{\mathbf{v}}|=5.0$, the impurity leaves behind a pronounced density wake and radiates excitations throughout the system above the critical velocity. In particular, the density depletion caused by the obstacle closely follows the obstacle when $\bm{\mathbf{v}}$ is small, while it lags far behind when $\bm{\mathbf{v}}$ is large.
[…]
@@ V Discussion and Outlook
The examples explored in Sec. IV demonstrate the breadth of physical phenomena accessible with this approach. These results establish PEPS–tVMC as a general-purpose tool for simulating elementary excitations, transport phenomena, and real-time response in a wide range of quantum systems with realistic computational resources. The improvements developed here apply equally to ground-state PEPS–VMC calculations [24, 25, 26, 27]. Extensions to excited states, finite temperature [76], and open quantum systems are also promising avenues. In the case of finite temperature, if purification-based ansatz is used, the small-$o$ trick can save the memory cost by a factor of 16 for spinful fermions.
@@ Acknowledgements.
Part of the algorithm is tested using the TenPy code base [77, 78]. The code for production is implemented with JAX [79]. I am supported by a startup grant from the IOP-CAS. I thank my colleagues Yu-Xiang Zhang, Shi-Xin Zhang, and Tao Xiang for sharing computational resources with me. I benefited greatly from discussion and collaboration on various related projects with Bartholomew Andrews, Zhehao Dai, Jannes Nys, Sheng-Hsuan Lin, Sajant Anand, Taige Wang, Michael Zaletel, Frank Pollmann, Wen-Yuan Liu, Jiahang Hu, Jiajun Yu, and Yuntian Gu. After the paper appeared as a preprint, I benefited from questions, comments, and discussion from Giuseppe Carleo, Markus Schmitt, Miles Stoudenmire, Joseph Tindall, Jad Halimeh, Qi Yang, and Yijia Wang. The data that support the findings of this article are openly available [80]; embargo periods may apply.
@@
    To implement this, one loops over $k$ and use an accumulator for the sum. This saves the memory by a factor of $K$ with no additional time cost. ($oo^{\dagger}$ is $K$-times faster than $OO^{\dagger}$, but one executes $K$ of them in sequence. The memory cost of $p_{sa}$ is negligible because its dtype is jnp.int8 compared to $O$’s dtype jnp.complex128).
@@ 5 OOdag : (Ns, Ns), obtained from get_OOdag()
  • dtdt: time step of the Runge-Kutta integrator
@@
S-3 Energy evolution
We use energy conservation to check the correctness of our simulations.
(a) Energy evolution of the Chern insulator. The relative energy drift is $2\times 10^{-4}$ over the evolution.
(b) Energy evolution of the bosonic Hofstadter model during charge pumping. The energy is not expected to be conserved during the pumping.
(c) Energy evolution of the $\mathbb{Z}_{2}$ lattice gauge theory in the deconfined phase. The relative energy drift is $1\times 10^{-3}$ over the evolution.
(d) Energy evolution of the $\mathbb{Z}_{2}$ lattice gauge theory in the Higgs phase. The relative energy drift is $2\times 10^{-3}$ over the evolution.
Figure S1: Energy evolutions of the simulations presented in the main text. In the case of bosonic Hofstadter model, because the Hamiltonian is explicitly time dependent, energy conservation is not expected and cannot be used as an accuracy diagnostic. Instead, we assess accuracy by comparing results obtained with two boundary-MPS bond dimensions, $D^{\prime}=24$ and $32$.